# Decentralized Collective World Model for Emergent Communication and Coordination

### v1

@@ Abstract
We propose a fully decentralized multi-agent world model that enables both symbol emergence for communication and coordinated behavior through temporal extension of collective predictive coding. Unlike previous research that focuses on either communication or coordination separately, our approach achieves both simultaneously. Our method integrates world models with communication channels, enabling agents to predict environmental dynamics, estimate states from partial observations, and share critical information through bidirectional message exchange with contrastive learning for message alignment. Using a two-agent trajectory drawing task, we demonstrate that our communication-based approach outperforms non-communicative models when agents have divergent perceptual capabilities, achieving the second-best coordination after centralized models. Importantly, our distributed approach with constraints preventing direct access to other agents’ internal states facilitates the emergence of more meaningful symbol systems that accurately reflect environmental states. These findings demonstrate the effectiveness of decentralized communication for supporting coordination while developing shared representations of the environment.
@@ I Introduction
Coordination through shared symbolic communication is fundamental to human society, enabling us to collectively achieve goals beyond individual capabilities [1]. In the real world, humans accomplish complex tasks through coordination based on commonly understood symbols and languages that emerge through distributed interactions. As environments increasingly integrate artificial systems with humans, a critical challenge remains unsolved: how to enable distributed multi-agent systems to simultaneously develop shared symbol systems and effective coordination without centralized control. This paper addresses this challenge by proposing an approach that integrates world models with communication channels, enabling agent groups to form symbol systems while coordinating through communication in partially observable environments. Our method allows agents to use world models to estimate environmental states from partial observations, share this information through emergent communication via bidirectional message exchange using the acquired common symbol system, and achieve coordination in complex, partially observable settings.
Coordination in multi-agent systems requires a shared symbol system where all participants interpret symbols consistently [2]. Communication plays a crucial role by allowing agents to exchange independently acquired information and infer environmental states [3], enabling them to complement each other’s knowledge and abilities, especially in partially observable environments where individual perception is limited [2, 4].
@@
World model-based approaches offer promising frameworks for coordination but currently lack mechanisms for collective knowledge formation. In partially observable environments, agents using world models can predict future states and infer hidden information [6, 7], potentially enhancing coordination capabilities [8]. However, current world model approaches in multi-agent settings primarily focus on individual agents’ learning rather than collective knowledge formation. Traditional multi-agent reinforcement learning methods often employ parameter sharing or learning signal transmission between agents, creating centralized systems [9, 10]. These systems become computationally expensive as the number of agents increases [11] and are impractical in real-world settings where centralized control is not feasible.
Decentralized approaches to multi-agent coordination lack mechanisms for developing shared symbol systems. Recent research has explored decentralized approaches that transmit information through non-differentiable messages without requiring parameter sharing or centralized learning [12, 13, 14], making them suitable for real-world applications where centralized control is impractical [15]. However, existing decentralized methods typically focus on immediate task achievement rather than developing shared symbol systems that capture the dynamic nature of environments. This limitation restricts their ability to form a common understanding that would enable more sophisticated coordination in complex, changing scenarios.
@@
This work makes three significant contributions to multi-agent coordination and communication research. First, we implement a multi-agent world model that can be learned in a fully decentralized manner based on the CPC framework. Second, we demonstrate that bidirectional message exchange using the acquired common symbol system promotes effective coordination behaviors in partially observable environments, allowing agents to complement each other’s limited perceptions. Third, we show that the distributed constraints of our approach lead to the formation of meaningful symbol systems that represent the global environmental state, emerging naturally from local observations. From the perspective of CPC research, our method extends previous CPC-based approaches temporally, providing a framework that captures contextual information. While our approach shares similarities with recent work by [26] in using contrastive learning for message alignment, we uniquely model agents as dynamic systems that predict environmental changes.
@@ III Proposed model
In this section, we first describe the model architecture and objective function. Next, we explain the approximation method for computing the collective term through symbol exchange. Finally, we detail the action determination procedure of agents within our proposed model.
Note that the primary contribution of this paper is the introduction of a multi-agent world model that facilitates symbol emergence for coordination. As elaborated in the Experiments section, both the world model and policies are trained using pre-generated expert data demonstrating cooperative behaviors. We do not address skill acquisition approaches such as reinforcement learning.
III-A CPC-based Decentralized Multi-agent World Model
Our proposed model, illustrated in Figure 2, is constructed around interconnected agent world models that exchange messages. Each world model learns environmental dynamics as state transitions in a latent space by modeling the observation generation process based on a POMDP. We employ the Recurrent State Space Model [29] as the foundational architecture for world models. Our key innovation is the introduction of a probabilistic variable representing messages exchanged between agents, which influences the generation of each agent’s latent variables at each time step. For a system with agents, the multi-agent world model comprises the following components:
Here, represent the observation, action, and latent variable of agent at time , respectively. The latent state encompasses both a probabilistic latent variable and a deterministic latent variable , where is the internal state of the Gated Recurrent Unit [30] in agent ’s world model. Additionally, denotes a probabilistic variable representing the common message shared between agents at time .
@@ III-B Approximation of the CR term for Distributed Learning
A significant challenge emerges in implementing the aforementioned model. While our theoretical formulation presupposes the existence of a model that infers messages from the latent variables of all agents, such a model cannot be directly computed in a decentralized system. In natural symbol emergence, an individual’s cognitive system is confined to its sensorimotor boundaries, rendering it impossible for an agent to directly access other agents’ internal representations. This constraint makes the CR term incalculable in its original form, presenting a substantial impediment to implementing distributed learning across agents.
To address this constraint, we propose an approximation methodology that enables independent learning for each agent. We introduce independent probabilistic variables representing each agent’s estimate of the shared message . We then approximate both the prior distribution and posterior distribution of the message using a Product-of-Experts (PoE) formulation:
@@
With this approximation established, we must determine how to set the prior distribution of messages for each agent. We propose defining as a distribution that reflects the collective prior knowledge at that time. Specifically, we consider a distribution based on what messages other agents are inferring at that moment. Thus, we define agent ’s prior distribution as a probability distribution represented by the PoE of other agents’ posterior distributions:
where denotes a normalization constant. This formulation establishes a feedback mechanism wherein each agent’s message inference is influenced by the collective understanding of all other agents, facilitating convergence to a shared communication protocol.
In this paper, we consider the case where the number of agents is . We designate the agent that infers messages from observations and transmits them to the other agent as the Speaker, and the agent that updates parameters based on observations and received messages from the other agent as the Listener. The Listener’s CR term at time is the KL divergence related to both agents’ message posterior distributions:
where represent the latent variables and messages of the Listener and Speaker, respectively. A practical challenge arises in calculating equation (III-B), as it requires the parameters of the message posterior distributions. However, in realistic scenarios, agents communicate by exchanging sampled messages rather than distributional parameters. Each agent samples messages from its posterior distribution and exchanges these samples with other agents. Consequently, an agent has access only to message samples from other agents, not to the parameters of their posterior distributions. To address this limitation, we leverage the Noise Contrastive Estimation framework [31] and employ the InfoNCE loss [32] to estimate the message inference model from samples alone:
To summarize our approach, we denote the agents as and . In practice, each agent learns by minimizing the weighted distributed CFE :
@@ IV-A Task Setup
We conducted experiments in a simulated environment to investigate whether multiple agents can achieve coordinated behavior through message formation that represents the complete environmental state. The experimental task involved two agents collaboratively moving a point in two-dimensional space to trace a predefined trajectory. Each agent receives sensory signals derived from the current coordinates of point in the x-y coordinate system, processed through agent-specific sensory modules. These modules discretize one coordinate axis into a predetermined number of bins and introduce noise. Specifically, agent A’s sensory module discretizes the y-axis coordinates, while agent B’s module discretizes the x-axis coordinates. By manipulating the number of bins in these sensory modules, we can systematically control the extent of environmental information accessible to each agent. With an infinite number of bins, no discretization occurs, enabling complete environmental perception. Conversely, with a single bin (bin=1), agents can observe only one axis, effectively reducing each agent’s perceptible environmental state by half. This configuration allows us to transition from a Markov Game setting (infinite bins) to a Decentralized Partially Observable Markov Decision Process (Dec-POMDP) setting (finite bins).

### v2

@@ Abstract
We propose a fully decentralized multi-agent world model that enables both symbol emergence for communication and coordinated behavior through temporal extension of collective predictive coding. Unlike previous research that focuses on either communication or coordination separately, our approach achieves both simultaneously. Our method integrates world models with communication channels, enabling agents to predict environmental dynamics, estimate states from partial observations, and share critical information through bidirectional message exchange with contrastive learning for message alignment. Using a two-agent trajectory drawing task, we demonstrate that our communication-based approach outperforms non-communicative models when agents have divergent perceptual capabilities, achieving the second-best coordination after centralized models. Importantly, our decentralized approach with constraints preventing direct access to other agents’ internal states facilitates the emergence of more meaningful symbol systems that accurately reflect environmental states. These findings demonstrate the effectiveness of decentralized communication for supporting coordination while developing shared representations of the environment.
@@ I Introduction
Coordination through shared symbolic communication is fundamental to human society, enabling us to collectively achieve goals beyond individual capabilities [1]. As environments increasingly integrate artificial systems with humans, a critical challenge remains unsolved: how to enable distributed multi-agent systems to simultaneously develop shared symbol systems and effective coordination without centralized control. This paper addresses this challenge by proposing an approach that integrates world models with communication channels, enabling agent groups to form symbol systems while coordinating through communication in partially observable environments. Our method allows agents to use world models to estimate environmental states from partial observations, share this information through emergent communication via bidirectional message exchange using the acquired common symbol system, and achieve coordination in complex, partially observable settings.
Coordination in multi-agent systems requires a shared symbol system where all participants interpret symbols consistently [2]. Communication plays a crucial role by allowing agents to exchange independently acquired information and infer environmental states [3], enabling them to complement each other’s knowledge and abilities, especially in partially observable environments [2, 4].
@@
World models-based approaches involve agents maintaining internal models that learn environmental dynamics and state estimation, forming latent state space representations from raw sensorimotor information [6, 7, 8]. These learned representations enable agents to predict future states and infer hidden environmental information from incomplete observations, potentially enhancing coordination capabilities in multi-agent settings [9]. However, current world model approaches in multi-agent settings lack mechanisms for collective knowledge formation through distributed symbol systems. Traditional multi-agent reinforcement learning methods often employ parameter sharing or centralized learning [10, 11], becoming computationally expensive and impractical in real-world settings [12].
Decentralized approaches to multi-agent coordination lack mechanisms for developing shared symbol systems. Recent research has explored decentralized approaches that transmit information through non-differentiable messages [13, 14, 15], making them suitable for real-world applications where centralized control is impractical [16]. However, existing decentralized methods typically focus on immediate task achievement rather than developing shared symbol systems that capture environmental dynamics. This limitation restricts their ability to form a common understanding that would enable more sophisticated coordination in complex, changing scenarios.
@@
This work makes three significant contributions to multi-agent coordination and communication research. First, we implement a two-agent world model that integrates CPC-based symbol emergence with temporal dynamics learning in a fully decentralized manner. Second, we demonstrate that environment-general symbols learned through world model prediction—rather than task-specific communication protocols—can effectively support coordination in partially observable environments, allowing agents to complement each other’s limited perceptions. Third, we show that the distributed constraints of our approach lead to the formation of meaningful symbol systems that represent the global environmental state, emerging naturally from predictive learning of environmental dynamics. From the perspective of CPC research, our method extends previous CPC-based approaches temporally, providing a framework that captures environmental dynamics rather than static observations. While our approach shares similarities with recent work by [27] in using contrastive learning for message alignment, we uniquely develop environment-general symbolic representations through predictive modeling.
@@ III Proposed model
In this section, we first describe the centralized model architecture and objective function for our two-agent system. Next, we explain how this centralized formulation is decomposed for decentralized implementation through approximation methods. Finally, we detail the action determination procedure of agents.
III-A CPC-based Multi-agent World Model
Following the CPC framework, we first formulate a centralized model that assumes access to shared messages between agents, then decompose it for decentralized implementation. Our proposed model, illustrated in Figure 2, is constructed around interconnected agent world models that jointly infer messages through exchange. Each world model learns environmental dynamics as state transitions in a latent space by modeling the observation generation process based on a POMDP. We employ the Recurrent State Space Model [30] as the foundational architecture for world models. Our key innovation is the introduction of a probabilistic variable representing messages exchanged between agents, which follows a continuous distribution enabling effectively infinite possible messages, and influences the generation of each agent’s latent variables at each time step. For a system with two agents (), the multi-agent world model comprises the following components:
Here, , and represent the observation, action, and latent variable of agent at time , respectively. The latent state encompasses both a probabilistic latent variable and a deterministic latent variable , where is the internal state of the Gated Recurrent Unit [31] in agent ’s world model. Additionally, denotes a probabilistic variable representing the common message shared between agents at time .
@@ III-B Approximation of the CR term for Distributed Learning
A significant challenge emerges in implementing the aforementioned model. While our theoretical formulation presupposes the existence of a model that infers messages from the latent variables of two agents, such a model cannot be directly computed in a decentralized system. In natural symbol emergence, an individual’s cognitive system is confined to its sensorimotor boundaries, rendering it impossible for an agent to directly access other agents’ internal representations. This constraint makes the CR term incalculable in its original form, presenting a substantial impediment to implementing distributed learning across agents.
To address this constraint, we propose an approximation methodology that enables independent learning for each agent. We introduce independent probabilistic variables and representing each agent’s estimate of the shared message . We then approximate both the prior distribution and posterior distribution of the message using a Product-of-Experts (PoE) formulation:
@@
With this approximation established, we must determine how to set the prior distribution of messages for each agent. We propose defining as a distribution that reflects the collective prior knowledge at that time. Specifically, we consider a distribution based on what messages the other agent is inferring at that moment. Thus, we define each agent’s prior distribution as the other agent’s posterior distribution:
This formulation establishes a feedback mechanism wherein each agent’s message inference is influenced by the collective understanding of the other agent, facilitating convergence to a shared communication protocol.
We designate the agent that infers messages from observations and transmits them to the other agent as the Speaker, and the agent that updates parameters based on observations and received messages from the other agent as the Listener. The Listener’s CR term at time is the KL divergence related to both agents’ message posterior distributions:
where and represent the latent variables and messages of the Listener and Speaker, respectively. A practical challenge arises in calculating equation (III-B), as it requires the parameters of the message posterior distributions. However, in realistic scenarios, agents communicate by exchanging sampled messages rather than distributional parameters. Each agent samples messages from its posterior distribution and exchanges these samples with other agents. Consequently, an agent has access only to message samples from other agents, not to the parameters of their posterior distributions. To address this limitation, we leverage the Noise Contrastive Estimation framework [32] and employ the InfoNCE loss [33] to estimate the message inference model from samples alone:
In practice, each agent learns by minimizing the weighted distributed CFE :
@@ IV-A Task Setup
We conducted experiments in a simulated environment to investigate whether multiple agents can achieve coordinated behavior through message formation that represents the complete environmental state. The experimental task involved two agents collaboratively moving a point in two-dimensional space to trace a predefined trajectory. Each agent receives sensory signals from point P coordinates through agent-specific sensory modules. These modules discretize one coordinate axis into a predetermined number of bins and introduce noise. Specifically, agent A’s sensory module discretizes the y-axis coordinates, while agent B’s module discretizes the x-axis coordinates. By manipulating the number of bins in these sensory modules, we can systematically control the extent of environmental information accessible to each agent. With an infinite number of bins, no discretization occurs, enabling complete environmental perception. Conversely, with a single bin (bin=1), agents can observe only one axis, effectively reducing each agent’s perceptible environmental state by half. This configuration allows us to transition from a Markov Game setting (infinite bins) to a Decentralized Partially Observable Markov Decision Process (Dec-POMDP) setting (finite bins).